I need to write as equation and solve.
One month's rent plus half of one months rent equals 1725
THanks

Answer by checkley77(12844)

X+X/2=1725
3X/2=1725
3X=2*1725
3X=3,450
X=3,450
X=$1,150 ANSWER FOR 1 MONTHS RENT.